# Nwāchā Munā: A Devanagari Speech Corpus and Proximal Transfer Benchmark for Nepal Bhasha ASR

### 08 Mar 2026 v1

@@ 4.2.  Results and Discussion
Strategies (model details CER (%)
Zero-Shot NepConformer 52.54
Semi-supervised NepConformer (Labelled + pseudo-labels, 14.72 hours) 19.83
NepConformer + 5-gram KenLM (Shallow fusion, $\lambda=0.4$, $\beta=1.5$, beam 128) 19.75
Decoder-only NepConformer (Encoder frozen, 5.39 hours) 18.77
Whisper-Small (Fine-tuned on base data, 5.39 hours) 18.76
NepConformer (Fine-tuned on base data, 5.39 hours) 18.72
Whisper-Small + Augmentation (Time stretch, pitch shift, Gaussian noise) 17.88
NepConformer + Augmented Data (23.05 hours) 17.59
Our experimental results demonstrate that transferring representations from a neighboring language with a similar script is a viable alternative to massive multilingual pre-training for low-resource ASR, suggesting that linguistic proximity can sometimes outweigh model scale in ultra-low-resource South Asian scenarios. Table 8 shows the orthographic and phonological comparison of the Nepali and Newari Language. The NepConformer model achieved a baseline CER of 18.72%, effectively matching the performance of the fine-tuned Whisper-Small model (18.76%) despite possessing significantly fewer parameters and lacking the huge data backing available in multilingual models like Whisper. This parity suggests that the acoustic and orthographic overlap between Nepali and Newari facilitates robust feature extraction, minimizing the need for the extensive capacity of large-scale models. Furthermore, the application of data augmentation proved critical in this data-scarce regime, yielding the state-of-the-art performance of 17.59% CER for NepConformer and 17.88% for Whisper-Small, highlighting that augmenting scarce audio data can effectively overcome dataset limitations and drive substantial performance gains.
Table 8: Orthological and Phonological Comparison of Nepali and Newari
@@
Finally, integrating the KenLM n-gram language model during decoding reduced the WER by approximately 11.7% relative to the baseline. At the same time, the CER increased slightly by 1.37%, as the language model tends to favor standard spellings, which can override phonetically correct but unusual or local forms in Newari speech. This happens partly because the text used to train the language model doesn’t always match how people actually speak. Overall, this shows the trade-off between making words more consistent and keeping the natural variations of this low-resource, agglutinative language.
@@ 4.3.  Error Analysis
Newari’s rich morphological structure, characterized by productive affixation and concatenated morphemes, makes character-level analysis particularly informative. Although the model frequently predicts individual characters correctly, it struggles to assemble them into accurate representations. As a result, the CER remains relatively low, while the WER is substantially higher, reflecting the difficulty of modeling agglutinative formations and complex orthographic concatenations (K et al., 2025).
@@
Table 10: Qualitative Error Analysis of NewaASR (underline in prediction column represents the error segment)
@@
If it is dirty, the shoes must be washed
@@ 5.  Conclusion
In this work, we developed resources to help bridge the digital divide for Nepal Bhasha (Newari) by curating a high-quality 5.39-hour transcribed speech corpus, with all transcripts encoded in the Devanagari script. Using this resource, we systematically evaluated cross-lingual transfer learning strategies, benchmarking NepConformer along with a multilingual Whisper-Small baseline. Our experiments demonstrate substantial gains from supervised fine-tuning and data augmentation, reducing the CER from a high zero-shot baseline to below 18%. The frozen-encoder fine-tuning strategy further enabled stable and effective adaptation in this ultra-low-resource setting. Moreover, shallow fusion with a KenLM $n$-gram model enhances lexical regularity and contributes to a reduction in WER, even when character-level improvements remain marginal, indicating improved word-level consistency in decoded outputs. Beyond Nepal Bhasa, our findings indicate that intra-regional transfer learning within South Asian language clusters may offer a computationally efficient pathway for scaling ASR to other endangered and minority languages
Ethics Statement and Limitations
This work offers meaningful societal and technological benefits for the Nepal Bhasha speaking community. By providing an open-source Devanagari speech corpus and establishing strong ASR benchmarks, the resources developed pave the way for voice-driven AI products that can significantly enhance digital accessibility. Furthermore, this research plays an important role in building extensive speech archives, allowing endangered language communities to actively document and preserve their linguistic heritage. Data collection was conducted respectfully within the community, engaging native speakers across different age groups and genders to ensure proper community representation.
Despite the promising benefits, this work faces several constraints related to ultra-low-resource settings. Primarily, the curated Nwāchā Munā dataset, while high-quality, remains relatively small compared to standard ASR benchmarks. Our current evaluation focuses predominantly on read speech and formal sentences, which may not fully capture the acoustic variability, disfluencies, and rapid speaking rates encountered in everyday spontaneous conversations. Also, external n-gram language model utilized during shallow fusion is constrained by domain mismatches and possesses limited coverage of natural, conversational linguistic forms.
Bibliographical References
@@
  • T. D. K, J. James, D. P. Gopinath, and M. A. K (2025) Advocating character error rate for multilingual ASR evaluation. In Findings of the Association for Computational Linguistics: NAACL 2025, L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 4941–4950. External Links: Link, Document, ISBN 979-8-89176-195-7 Cited by: §4.3.
  • M. Meelen, A. O'Neill, and R. Coto-Solano (2024) End-to-end speech recognition for endangered languages of nepal. In Proceedings of the 7th Workshop on Computational Methods for Endangered Languages (ComputEL-7), pp. 83–93. External Links: Link Cited by: §2, §3.1.2.
@@
  • National Statistics Office (2023) National population and housing census 2021: national report on caste/ethnicity, language & religion. Government of Nepal, Kathmandu, Nepal. External Links: Link Cited by: §3.1.2.

### 30 Mar 2026 v2

@@ 4.2.   Findings and Analysis
Strategies (model details) Dataset CER (%) WER (%)
Zero-Shot NepConformer - 52.54 98.68
Semi-supervised NepConformer 5.39 hours labelled
+ 9.33 hours Pseudo labeled
19.83 58.28
NepConformer + 5-gram KenLM (Shallow fusion, $\lambda=0.4$, $\beta=1.5$, beam 128) 5.39 hours labelled 19.75 48.46
Decoder-only NepConformer (Encoder frozen) 5.39 hours labelled 18.77 53.47
Whisper-Small (Fine-tuned on base data) 5.39 hours labelled 18.76 51.44
NepConformer (Fine-tuned on base data) 5.39 hours labelled 18.72 54.34
Whisper-Small + Augmentation (Time stretch, pitch shift, Gaussian noise) 5.39 hours labelled 17.88 48.80
NepConformer + Augmented Data (speed perturbation ($0.9\times,1.1\times$), volume randomization, and noise injection) 23.05 hours labelled 17.59 52.61
Due to differences in script representation, normalization, evaluation protocol and the unavailability of their exact test split, a direct numerical comparison with Meelen et al. (2024) is not feasible. Instead, our experimental results demonstrate that transferring representations from a neighboring language with a similar script is a viable alternative to massive multilingual pre-training for low-resource ASR, suggesting that linguistic proximity can potentially outweigh model scale in ultra-low-resource South Asian scenarios.
Table 9 shows the orthographic and phonological comparison of the Nepali and Newari Language. Table 10 presents a comparison of model size and computational resource requirements, highlighting that NepConformer is significantly more lightweight, with only 30.54 million parameters and lower VRAM usage (8.56 GB) compared to Whisper-Small, which requires substantially higher usage (12.38 GB).
Despite possessing significantly fewer parameters and lacking the huge data backing available in multilingual models like Whisper, the NepConformer model achieved a baseline CER of 18.72%, effectively matching the performance of the fine-tuned Whisper-Small model (18.76%). This parity suggests that the acoustic and orthographic overlap between Nepali and Newari facilitates robust feature extraction, minimizing the need for the extensive capacity of large-scale models. Furthermore, the application of data augmentation proved critical in this data-scarce regime, yielding the state-of-the-art performance of 17.59% CER for NepConformer and 17.88% for Whisper-Small, highlighting that augmenting scarce audio data can effectively overcome dataset limitations and drive substantial performance gains.
Table 9: Orthographic and Phonological Comparison of Nepali and Newari
@@
Table 10: Computational Resource and Parameter Comparison
Model Parameter (Million) VRAM Used (GB)
NepConformer 30.54 8.56
Whisper(small) 244 12.38
@@
Additionally, the integration of the KenLM n-gram language model was evaluated across all models during decoding. The best results were observed with the NepConformer fine-tuned model, where WER was reduced by approximately 11.7% relative to the baseline. At the same time, the CER increased slightly by 1.37%, as the language model tends to favor standard spellings, which can override phonetically correct but unusual or local forms in Newari speech. This happens partly because the text used to train the language model doesn’t always match how people actually speak. Overall, this shows the trade-off between making words more consistent and keeping the natural variations of this low-resource, agglutinative language.
Finally, in the semi-supervised learning approach, incorporating the pseudo-labeled data with the original labeled training set degraded performance, increasing the CER to 19.83% compared to the 17.59% baseline. A qualitative analysis of a random sample of these pseudo-labels revealed three primary sources of error: extensive code-mixing (alternating with Nepali or English), frequent overlapping conversational speech, and highly variable acoustic environments (e.g., background music, distant microphones). Ultimately, this decline highlights the critical challenge that the severe domain shift introduced by uncurated data can completely overshadow the benefits of increased training volume. This suggests that strict domain alignment, speaker diarization, and rigorous filtering are significantly more crucial than raw data quantity for endangered-language ASR.
@@ 4.3.  Error Analysis
Newari’s rich morphological structure, characterized by productive affixation and concatenated morphemes, makes character-level analysis particularly informative. Although the model frequently predicts individual characters correctly, it struggles to assemble them into accurate representations. As a result, the CER remains relatively low, while the WER is substantially higher, reflecting the difficulty of modeling agglutinative formations and complex orthographic concatenations. This is also observed in other work (D K et al., 2025).
@@
Table 12: Qualitative Error Analysis of Nwāchā Munā (underline in prediction column represents the error segment)
@@
If the shoe is dirty, it must be washed
@@ 5.  Conclusion
In this work, we developed resources to help bridge the digital divide for Nepal Bhasha (Newari) by curating a high-quality 5.39-hour transcribed speech corpus, with all transcripts encoded in the Devanagari script. Using this resource, we systematically evaluated cross-lingual transfer learning strategies, benchmarking NepConformer along with a multilingual Whisper-Small baseline. Our experiments demonstrate substantial gains from supervised fine-tuning and data augmentation, reducing the CER from a high zero-shot baseline to below 18%. The frozen-encoder fine-tuning strategy further enabled stable and effective adaptation in this ultra-low-resource setting. Moreover, shallow fusion with a KenLM $n$-gram language model enhances lexical regularity and contributes to a reduction in WER, even when character-level improvements remain marginal, indicating improved word-level consistency in decoded outputs. Beyond Nepal Bhasha, our findings indicate that intra-regional transfer learning within South Asian language clusters may offer a computationally efficient pathway for scaling ASR to other endangered and minority languages
Data and Code Availability
The training scripts and source code are available on GitHub (https://github.com/ilprl/nwacha-muna). The model weights and audio datasets are hosted on Hugging Face (https://huggingface.co/collections/ilprl-docse/nwacha-muna).
6.  Ethics Statement and Limitations
This work offers meaningful societal and technological benefits for the Nepal Bhasha speaking community. By providing an open-source Devanagari speech corpus and establishing strong ASR benchmarks, the resources developed pave the way for voice-driven AI products that can significantly enhance digital accessibility. Furthermore, this research plays an important role in building extensive speech archives, allowing endangered language communities to actively document and preserve their linguistic heritage.
The dataset was collected through voluntary participation from the community, involving native speakers across diverse age groups and genders to ensure representative coverage. For the recorded audio, consent was obtained from all volunteers, and no sensitive or personally identifiable information (PII) was intentionally included in the spoken transcripts. Nonetheless, we acknowledge that models trained on limited data may reflect embedded societal biases or result in acoustic-phonetic misrepresentations, particularly in ultra-low-resource settings. Because of these limitations, the baseline ASR system developed in this study should not be deployed in critical applications without rigorous human verification. The primary objective of this work remains the promotion of linguistic inclusion and digital support for under-resourced language communities.
Despite the promising benefits, this work faces several constraints related to ultra-low-resource settings. Primarily, the curated Nwāchā Munā dataset, while high-quality, remains relatively small compared to standard ASR benchmarks. Our current evaluation focuses predominantly on read speech and formal sentences, which may not fully capture the acoustic variability, disfluencies, and rapid speaking rates encountered in everyday spontaneous conversations. Also, external n-gram language model utilized during shallow fusion is constrained by domain mismatches and possesses limited coverage of natural, conversational linguistic patterns.
7.  Bibliographical References
@@
  • T. D K, J. James, D. P. Gopinath, and M. A. K (2025) Advocating character error rate for multilingual ASR evaluation. In Findings of the Association for Computational Linguistics: NAACL 2025, L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 4941–4950. External Links: Link, Document, ISBN 979-8-89176-195-7 Cited by: §4.3.
@@
  • D. M. Eberhard, G. F. Simons, and C. D. Fennig (2026) Ethnologue: languages of the world. Twenty-ninth edition, SIL International, Dallas, Texas. Note: Online version: http://www.ethnologue.com Cited by: §1.
@@
  • M. Meelen, A. O'Neill, and R. Coto-Solano (2024) End-to-end speech recognition for endangered languages of nepal. In Proceedings of the 7th Workshop on Computational Methods for Endangered Languages (ComputEL-7), pp. 83–93. External Links: Link Cited by: §2, §3.1.2, §4.2.
@@
  • K. Pantha, R. R. Ghimire, and B. K. Bal (2025) Speech personalization using parameter efficient fine-tuning for Nepali speakers. In Proceedings of the 5th Conference on Language, Data and Knowledge: Fifth Workshop on Language Technology for Equality, Diversity, Inclusion, K. Gkirtzou, S. Žitnik, J. Gracia, D. Gromann, M. P. di Buono, J. Monti, and M. Ionov (Eds.), Naples, Italy, pp. 190–199. External Links: Link, ISBN 978-88-6719-334-9 Cited by: §2.
@@
  • S. Paudel, B. K. Bal, and D. Shrestha (2023) Large vocabulary continous speech recognition for Nepali language using CNN and transformer. In Proceedings of the 4th Conference on Language, Data and Knowledge, S. Carvalho, A. F. Khan, A. O. Anić, B. Spahiu, J. Gracia, J. P. McCrae, D. Gromann, B. Heinisch, and A. Salgado (Eds.), Vienna, Austria, pp. 328–333. External Links: Link Cited by: §2.
@@
  • S. Regmi and B. K. Bal (2021) An end-to-end speech recognition for the Nepali language. In Proceedings of the 18th International Conference on Natural Language Processing (ICON), S. Bandyopadhyay, S. L. Devi, and P. Bhattacharyya (Eds.), National Institute of Technology Silchar, Silchar, India, pp. 180–185. External Links: Link Cited by: §2.